# SlumpGuard: An AI-Powered Real-Time System for Automated Concrete Slump Prediction via Video Analysis

### v3

@@ 1 Introduction
Concrete stands as the backbone of modern construction infrastructure, with over 14 billion cubic meters produced annually worldwide, making it the most widely consumed material after water [hoang2016:estimating]. Among the critical properties governing concrete performance, workability, which is primarily assessed through slump testing, represents a fundamental parameter that directly influences mixing efficiency, transportation logistics, placement operations, and compaction effectiveness on construction sites [chen2022:research, ojala2024:estimating]. However, the complex and dynamic nature of construction environments presents unprecedented challenges for maintaining consistent concrete quality control. In particular, conventional manual slump testing, still the industry standard, often fails to meet the demanding requirements of large-scale, time-sensitive projects.
@@
The construction industry therefore continues to struggle with real-time quality assurance. On-site inspection of fresh concrete is typically performed only intermittently rather than for every delivered batch. For example, the Korean Construction Specification KCS 14 20 10 mandates slump inspection only once per day or once every 120 cubic meters (approximately 20 truckloads), making continuous, batch-level verification infeasible in practice. As a result, quality deviations that arise during mixing, transportation, or placement may remain undetected until concrete is already being poured, leading to delays, increases in cost, and inconsistency due to human measurement variability [tuan2021:situ].
@@
To address these challenges, researchers have increasingly explored automated and data-driven methods for assessing fresh concrete properties. As part of these efforts, sensor-based approaches employ piezoelectric, depth, and fiber-optic measurements to quantify workability-related characteristics. Leveraging these sensing modalities, recent work has explored combining the captured signals with deep learning and IoT-based frameworks to enable automated analysis of fresh concrete properties [han2025:field, kim2018:visualization, govindaraju2025:real]. However, sensor-based solutions heavily rely on hardware installation and calibration, and their practicality is often constrained by on-site regulatory restrictions and variations in equipment conditions across different construction environments.
@@
To overcome these limitations, visual information–based methods for slump estimation have also been proposed. Yucelai et al. [yucelai:ai] introduced a stereo-vision system for predicting slump during laboratory slump tests. Other studies attempted to estimate slump by installing cameras in proximity to the concrete mixer and analyzing captured images [idrees2024:automatic, hao2025:concrete]. Despite their promise, these approaches face significant barriers to real-world deployment. The method of Yucelai et al. [yucelai:ai] requires repeated on-site slump testing, which is rarely performed during routine concrete placement, limiting its practical applicability. Methods based on mixer-mounted cameras [idrees2024:automatic, hao2025:concrete] are also susceptible to dust and slurry contamination, leading to frequent maintenance and reduced operational reliability. Moreover, image-only frameworks generally fail to capture the flow-driven characteristics of fresh concrete, making it difficult to reflect the inherent dynamic behavior underlying slump.
@@ 2.1 Optical Flow
Optical flow is a computer vision technique used to estimate the motion of objects or scenes across video frames by analyzing the movement of pixels. It represents motion as a vector field, where each vector indicates the displacement of a pixel between consecutive frames. There are two main approaches: dense optical flow, which computes motion for every pixel, and sparse optical flow, which estimates motion only at selected feature points. While dense optical flow [horn1981:determining] offers a detailed view of the entire scene’s motion, it is computationally intensive. Sparse optical flow is more efficient but may overlook fine-grained motion.
In this work, we adopt the Lucas-Kanade method [lucas1981:iterative], a widely used algorithm for sparse optical flow, suitable for real-time applications such as tracking concrete flow. It assumes that the intensity of a point remains constant over time:
@@
FlowNet [dosovitskiy2015:flownet] pioneered the use of convolutional neural networks for end-to-end optical flow prediction, learning motion patterns directly from synthetic datasets. Building on this idea, FlowNet2 [ilg2017:flownet], PWC-Net [sun2018:pwc], and LiteFlowNet [hui2018:liteflownet] introduced architectural improvements such as pyramid-based warping, cost volume reasoning, and hierarchical refinement. More recently, RAFT [teed2020:raft] achieved significant performance gains by leveraging dense, all-pairs correlations and iterative updates to produce highly accurate flow estimates.
Previous studies have applied optical flow combined with imaging techniques for slump prediction under controlled experimental settings [tuan2021:situ, kim2018:visualization]. In contrast, our approach is specifically designed for robust, real-world deployment on construction sites, enabling continuous, automated slump quality monitoring under diverse and dynamic field conditions.
@@ 2.2 Object Detection
Object detection is a fundamental computer vision task widely applied in construction to automatically identify and localize key entities such as workers, equipment, and materials in images or videos. Recent studies have demonstrated its effectiveness in construction site monitoring [pfitzner2025:monitoring, kim2023:small].
State-of-the-art detection frameworks such as Faster R-CNN [ren2015:faster], Mask R-CNN [he2017:mask], SSD [liu2016:ssd], and YOLO [redmon2016:you] offer a good balance between speed and accuracy, making them well-suited for real-time applications on construction sites.
However, in construction environments where objects like materials and equipment often appear at arbitrary angles, standard axis-aligned bounding boxes may be insufficient. In such cases, oriented object detection becomes essential. Early approaches like R2CNN [jiang2017:r2cnn] introduced rotation-invariant proposals, while more recent models such as Oriented R-CNN [xie2021:oriented] and ReDet [han2021:redet] employ rotation-aware architectures to improve detection accuracy. These methods are particularly effective for automated inspection tasks involving rotated or non-axis-aligned objects.
@@ 2.3 Video Classification
Video classification is a fundamental task in computer vision that involves assigning semantic labels to video clips by analyzing both spatial and temporal information inherent in the data. Recent advances in this domain have been largely driven by two primary classes of deep learning architectures: 3D convolutional neural networks (3D CNNs) and vision transformers [dosovitskiy2020:image] tailored for video.
3D CNNs like ResNet3D [tran2018:closer] extend 2D convolutions to the temporal dimension, effectively capturing spatiotemporal features and achieving strong results on benchmarks such as Kinetics-400 [kay2017:kinetics]. Transformer-based models, notably TimeSformer [bertasius2021:space], use divided spatial and temporal self-attention to model long-range dependencies efficiently, surpassing many convolutional approaches. Further advancements include VideoSwin [liu2022:video] and MViT [fan2021:multiscale], which incorporate hierarchical and multiscale attention mechanisms to improve video representation learning and classification performance.
These video classification models have also been increasingly applied in the construction industry, playing a pivotal role in streamlining quality supervision, activity recognition, and concrete workability assessment. Previous studies have also explored video-based approaches for material quality monitoring and evaluation [guo2025:egocentric, ojala2024:estimating, idrees2024:automatic], alongside behavior classification systems aimed at improving worker safety and productivity [li2022:action, yang2023:transformer].
@@ 4.1 Chute Detection
To automatically identify when concrete is falling from the chute, we first detect the chute region in the video using an object detection model (see Step 1 in Figure 3). We employ YOLOv8 [yolov8], a state-of-the-art object detection model, known for its fast inference performance and high accuracy, making it suitable for real-time object detection. We initialize the model with weights pretrained on the DOTA dataset [xia2018:dota], which is designed for detecting oriented objects in aerial imagery and is well-suited for our task of detecting rotated chutes. Using this model, we simultaneously predict both axis-aligned and rotated-bounding boxes for the chutes, enabling us to detect the chute area in pouring videos. The detected chute is represented by a bounding box with parameters $(x,y,w,h,\theta)$, indicating the center coordinates, width, height, and rotation angle of the box.
@@ 4.3 Slump Prediction and Anomaly Detection
To enable efficient and real-time estimation of concrete slump, we leverage the drop timing and location identified in Section 4.2. Specifically, given the drop time $t$, we predict the slump using video data from the corresponding chute region. This process corresponds to Step 3 of the overall pipeline shown in Figure 3. To focus on the continuous flow of concrete, we adopt a video classification model, ResNet 3D [tran2018:closer]. Let $V$ denote the chute video (i.e., $V^{R}$ or $V^{L}$), and $t$ the detected drop time.
@@ 4.3.2 Data Augmentation
Given that construction sites are exposed to a wide range of lighting and environmental conditions, data augmentation is essential to improve the model’s generalization performance. To simulate these real-world variations, we employ a set of augmentation techniques, including ColorJitter, gamma correction, horizontal flipping, and contrast adjustment. We also apply MixUp [zhang2017:mixup] and CutMix [yun2019:cutmix], which blend either the pixel content or the labels of two training examples to generate mixed samples. While MixUp interpolates both images and labels, CutMix replaces a region of an image with a patch from another image, adjusting the labels proportionally to the area. These strategies encourage the model to learn more robust and smoother decision boundaries, and to be less sensitive to spurious correlations between local image regions and class labels.
@@ 4.3.3 Model
We employed the ResNet 3D architecture [tran2018:closer] as our video classification backbone, designed to jointly capture spatial and temporal patterns. We experimented with two variants: Mixed Convolution and (2+1)D convolution. The Mixed Convolution model uses 3D convolutions in early layers for motion modeling and 2D convolutions in later layers for efficient spatial reasoning, achieving performance comparable to full 3D models with fewer parameters. The (2+1)D model factorizes 3D convolutions into separate spatial and temporal steps, enhancing optimization and model capacity.
@@ 4.3.4 Training
To formulate the slump estimation task as a classification problem, we categorized the continuous slump values into five categorical intervals (e.g. 150–180mm, 180–210mm), where the width of each interval reflects the $\pm 30$ mm tolerance specified in the KS F 4009 standard [ksf4009], along with an additional margin for human error. For intervals at the extremes (i.e. below 150 mm and above 240 mm), which are rarely encountered in practice, we did not apply explicit error margins.
@@ 5.4.1 Model Training Setup
The model was optimized using the AdamW optimizer [loshchilov2017:decoupled] with a learning rate of 1e41e-4, and a weight decay of 1e41e-4. We used a total batch size of 128 and trained for 10 epochs using the OneCycle learning rate scheduler [smith2019:super] with cosine annealing, with a peak learning rate set to 1e31e-3.
@@
Transformer TimeSFormer [bertasius2021:space] 0.6980 0.7280 0.5724 0.6155
@@ 5.4.2 Results
As shown in Table 3, our models employing ResNet-3D as the backbone achieve over 80% accuracy on both the validation and test sets. In contrast, the Transformer [vaswani2017:attention]-based TimeSFormer [bertasius2021:space] exhibits lower accuracy. Among the evaluated architectures, we selected R(2+1)D-18 as our final model, as it achieved the highest performance on both validation and test sets. Based on this architecture, we conducted a detailed ablation study, and the corresponding results are reported in Table 3.

### v4

@@ 1 Introduction
Concrete stands as the backbone of modern construction infrastructure, with over 14 billion cubic meters produced annually worldwide, making it the most widely consumed material after water [12]. Among the critical properties governing concrete performance, workability, which is primarily assessed through slump testing, represents a fundamental parameter that directly influences mixing efficiency, transportation logistics, placement operations, and compaction effectiveness on construction sites [2, 28]. However, the complex and dynamic nature of construction environments presents unprecedented challenges for maintaining consistent concrete quality control. In particular, conventional manual slump testing, still the industry standard, often fails to meet the demanding requirements of large-scale, time-sensitive projects.
@@
The construction industry therefore continues to struggle with real-time quality assurance. On-site inspection of fresh concrete is typically performed only intermittently rather than for every delivered batch. For example, the Korean Construction Specification KCS 14 20 10 mandates slump inspection only once per day or once every 120 cubic meters (approximately 20 truckloads), making continuous, batch-level verification infeasible in practice. As a result, quality deviations that arise during mixing, transportation, or placement may remain undetected until concrete is already being poured, leading to delays, increases in cost, and inconsistency due to human measurement variability [36].
@@
To address these challenges, researchers have increasingly explored automated and data-driven methods for assessing fresh concrete properties. As part of these efforts, sensor-based approaches employ piezoelectric, depth, and fiber-optic measurements to quantify workability-related characteristics. Leveraging these sensing modalities, recent work has explored combining the captured signals with deep learning and IoT-based frameworks to enable automated analysis of fresh concrete properties [8, 20, 6]. However, sensor-based solutions heavily rely on hardware installation and calibration, and their practicality is often constrained by on-site regulatory restrictions and variations in equipment conditions across different construction environments.
@@
To overcome these limitations, visual information–based methods for slump estimation have also been proposed. Yucelai et al. [41] introduced a stereo-vision system for predicting slump during laboratory slump tests. Other studies attempted to estimate slump by installing cameras in proximity to the concrete mixer and analyzing captured images [15, 10]. Despite their promise, these approaches face significant barriers to real-world deployment. The method of Yucelai et al. [41] requires repeated on-site slump testing, which is rarely performed during routine concrete placement, limiting its practical applicability. Methods based on mixer-mounted cameras [15, 10] are also susceptible to dust and slurry contamination, leading to frequent maintenance and reduced operational reliability. Moreover, image-only frameworks generally fail to capture the flow-driven characteristics of fresh concrete, making it difficult to reflect the inherent dynamic behavior underlying slump.
@@ 2.1 Optical Flow
Optical flow is a computer vision technique used to estimate the motion of objects or scenes across video frames by analyzing the movement of pixels. It represents motion as a vector field, where each vector indicates the displacement of a pixel between consecutive frames. There are two main approaches: dense optical flow, which computes motion for every pixel, and sparse optical flow, which estimates motion only at selected feature points. While dense optical flow [13] offers a detailed view of the entire scene’s motion, it is computationally intensive. Sparse optical flow is more efficient but may overlook fine-grained motion.
In this work, we adopt the Lucas-Kanade method [27], a widely used algorithm for sparse optical flow, suitable for real-time applications such as tracking concrete flow. It assumes that the intensity of a point remains constant over time:
@@
FlowNet [4] pioneered the use of convolutional neural networks for end-to-end optical flow prediction, learning motion patterns directly from synthetic datasets. Building on this idea, FlowNet2 [16], PWC-Net [33], and LiteFlowNet [14] introduced architectural improvements such as pyramid-based warping, cost volume reasoning, and hierarchical refinement. More recently, RAFT [34] achieved significant performance gains by leveraging dense, all-pairs correlations and iterative updates to produce highly accurate flow estimates.
Previous studies have applied optical flow combined with imaging techniques for slump prediction under controlled experimental settings [36, 20]. In contrast, our approach is specifically designed for robust, real-world deployment on construction sites, enabling continuous, automated slump quality monitoring under diverse and dynamic field conditions.
@@ 2.2 Object Detection
Object detection is a fundamental computer vision task widely applied in construction to automatically identify and localize key entities such as workers, equipment, and materials in images or videos. Recent studies have demonstrated its effectiveness in construction site monitoring [29, 21].
State-of-the-art detection frameworks such as Faster R-CNN [31], Mask R-CNN [11], SSD [24], and YOLO [30] offer a good balance between speed and accuracy, making them well-suited for real-time applications on construction sites.
However, in construction environments where objects like materials and equipment often appear at arbitrary angles, standard axis-aligned bounding boxes may be insufficient. In such cases, oriented object detection becomes essential. Early approaches like R2CNN [17] introduced rotation-invariant proposals, while more recent models such as Oriented R-CNN [39] and ReDet [9] employ rotation-aware architectures to improve detection accuracy. These methods are particularly effective for automated inspection tasks involving rotated or non-axis-aligned objects.
@@ 2.3 Video Classification
Video classification is a fundamental task in computer vision that involves assigning semantic labels to video clips by analyzing both spatial and temporal information inherent in the data. Recent advances in this domain have been largely driven by two primary classes of deep learning architectures: 3D convolutional neural networks (3D CNNs) and vision transformers [3] tailored for video.
3D CNNs like ResNet3D [35] extend 2D convolutions to the temporal dimension, effectively capturing spatiotemporal features and achieving strong results on benchmarks such as Kinetics-400 [19]. Transformer-based models, notably TimeSformer [1], use divided spatial and temporal self-attention to model long-range dependencies efficiently, surpassing many convolutional approaches. Further advancements include VideoSwin [25] and MViT [5], which incorporate hierarchical and multiscale attention mechanisms to improve video representation learning and classification performance.
These video classification models have also been increasingly applied in the construction industry, playing a pivotal role in streamlining quality supervision, activity recognition, and concrete workability assessment. Previous studies have also explored video-based approaches for material quality monitoring and evaluation [7, 28, 15], alongside behavior classification systems aimed at improving worker safety and productivity [23, 40].
@@ 4.1 Chute Detection
To automatically identify when concrete is falling from the chute, we first detect the chute region in the video using an object detection model (see Step 1 in Figure 3). We employ YOLOv8 [18], a state-of-the-art object detection model, known for its fast inference performance and high accuracy, making it suitable for real-time object detection. We initialize the model with weights pretrained on the DOTA dataset [38], which is designed for detecting oriented objects in aerial imagery and is well-suited for our task of detecting rotated chutes. Using this model, we simultaneously predict both axis-aligned and rotated-bounding boxes for the chutes, enabling us to detect the chute area in pouring videos. The detected chute is represented by a bounding box with parameters $(x,y,w,h,\theta)$, indicating the center coordinates, width, height, and rotation angle of the box.
@@ 4.3 Slump Prediction and Anomaly Detection
To enable efficient and real-time estimation of concrete slump, we leverage the drop timing and location identified in Section 4.2. Specifically, given the drop time $t$, we predict the slump using video data from the corresponding chute region. This process corresponds to Step 3 of the overall pipeline shown in Figure 3. To focus on the continuous flow of concrete, we adopt a video classification model, ResNet 3D [35]. Let $V$ denote the chute video (i.e., $V^{R}$ or $V^{L}$), and $t$ the detected drop time.
@@ 4.3.2 Data Augmentation
Given that construction sites are exposed to a wide range of lighting and environmental conditions, data augmentation is essential to improve the model’s generalization performance. To simulate these real-world variations, we employ a set of augmentation techniques, including ColorJitter, gamma correction, horizontal flipping, and contrast adjustment. We also apply MixUp [43] and CutMix [42], which blend either the pixel content or the labels of two training examples to generate mixed samples. While MixUp interpolates both images and labels, CutMix replaces a region of an image with a patch from another image, adjusting the labels proportionally to the area. These strategies encourage the model to learn more robust and smoother decision boundaries, and to be less sensitive to spurious correlations between local image regions and class labels.
@@ 4.3.3 Model
We employed the ResNet 3D architecture [35] as our video classification backbone, designed to jointly capture spatial and temporal patterns. We experimented with two variants: Mixed Convolution and (2+1)D convolution. The Mixed Convolution model uses 3D convolutions in early layers for motion modeling and 2D convolutions in later layers for efficient spatial reasoning, achieving performance comparable to full 3D models with fewer parameters. The (2+1)D model factorizes 3D convolutions into separate spatial and temporal steps, enhancing optimization and model capacity.
@@ 4.3.4 Training
To formulate the slump estimation task as a classification problem, we categorized the continuous slump values into five categorical intervals (e.g. 150–180mm, 180–210mm), where the width of each interval reflects the $\pm 30$ mm tolerance specified in the KS F 4009 standard [22], along with an additional margin for human error. For intervals at the extremes (i.e. below 150 mm and above 240 mm), which are rarely encountered in practice, we did not apply explicit error margins.
@@ 5.4.1 Model Training Setup
The model was optimized using the AdamW optimizer [26] with a learning rate of 1e41e-4, and a weight decay of 1e41e-4. We used a total batch size of 128 and trained for 10 epochs using the OneCycle learning rate scheduler [32] with cosine annealing, with a peak learning rate set to 1e31e-3.
@@
Transformer TimeSFormer [1] 0.6980 0.7280 0.5724 0.6155
@@ 5.4.2 Results
As shown in Table 3, our models employing ResNet-3D as the backbone achieve over 80% accuracy on both the validation and test sets. In contrast, the Transformer [37]-based TimeSFormer [1] exhibits lower accuracy. Among the evaluated architectures, we selected R(2+1)D-18 as our final model, as it achieved the highest performance on both validation and test sets. Based on this architecture, we conducted a detailed ablation study, and the corresponding results are reported in Table 3.